# Probabilistic-based Feature Embedding of 4-D Light Fields for Compressive Imaging and Denoising

### v1

@@ Abstract
The high-dimensional nature of the 4-D light field (LF) poses great challenges in efficient and effective feature embedding that severely impact the performance of downstream tasks. To tackle this crucial issue, in contrast to existing methods with empirically-designed architectures, we propose probabilistic-based feature embedding (PFE), which learns a feature embedding architecture by assembling various low-dimensional convolution patterns in a probability space for fully capturing spatial-angular information. Building upon the proposed PFE, we then leverage the intrinsic linear imaging model of the coded aperture camera to construct a cycle-consistent 4-D LF reconstruction network from coded measurements. Moreover, we incorporate PFE into an iterative optimization framework for 4-D LF denoising. Our extensive experiments demonstrate the significant superiority of our methods on both real-world and synthetic 4-D LF images, both quantitatively and qualitatively, when compared with state-of-the-art methods. The source code will be publicly available at https://github.com/lyuxianqiang/LFCA-CR-NET.
@@ 1 Introduction
To process the 4-D LF, previous learning-based methods simply apply 2-D or 3-D convolutional filters to the stack of sub-aperture images (SAIs) Inagaki et al. (2018); Vadathya et al. (2019); Shin et al. (2018); Zhang et al. (2019), which has a limited ability to explore the angular relations among SAIs. Some advanced network architectures have been designed to model the LF structure better, such as the 4-D convolutional layer Yeung et al. (2018a), the epipolar plane images (EPIs) convolutional layer Wu et al. (2017, 2019); Heber et al. (2017), the spatial-angular separable (SAS) convolutional layer Guo et al. (2020); Yeung et al. (2018b); Jin et al. (2020a, b), and the spatial-angular interactive network Wang et al. (2020a). However, as these methods combine the spatial and angular features of the LF data in an intuitive or empirical manner, the optimal strategy to model the LF data still needs to be studied.
@@
In this paper, to embed 4-D LF data efficiently and effectively, we propose a probabilistic-based feature embedding module. As shown in Fig. 1, we formulate the problem in a probability space and propose to approximate a maximum posterior distribution (MAP) of a set of carefully-defined LF processing events, including both layer-wise spatial-angular feature extraction and network-level feature aggregation. Through droppath from a densely-connected template network, we derive an adaptively probabilistic-based feature embedding, which is sharply contrasted with existing manners that combine spatial and angular features empirically. Building upon the proposed probabilistic-based feature embedding method, we propose two distinct frameworks for the essential tasks of compressive LF imaging and LF image denoising. These tasks are critical in the context of LF data transmission and preprocessing for subsequent applications.
@@
Earlier 4-D LF capturing devices, such as camera array Wilburn et al. (2005) and camera gantry Vaibhav and Andrew , are bulky and costly to capture dense LFs. Although more recent commercial LF cameras, such as Lytro (2016) and RayTrix , are more convenient and efficient, the limited sensor resolution results in a trade-off between spatial and angular resolution. Besides, it is challenging to transmit and store such a large amount of 4-D LF data. The coded aperture camera can encode the 4-D LF to 2-D coded measurements (CMs) without losing the spatial resolution, which is a promising way for 4-D LF acquisition, transmission, and storage. The bottleneck lies in the ability of the subsequent reconstruction algorithm, which reconstructs the 4-D LF image from the 2-D CMs. Although existing deep learning-based 4-D LF reconstruction methods from CMs Inagaki et al. (2018); Vadathya et al. (2019) have presented significantly better reconstruction quality than conventional ones Babacan et al. (2012); Marwah et al. (2013), these methods employ plain convolutional networks that are purely data-driven without taking the observation model into account. More recently, Guo et al. (2020) proposed an unrolling-based method to link the observation model of coded apertures and deep learning elegantly and improve the LF reconstruction quality significantly. In this paper, we propose a physically interpretable framework incorporated with the probabilistic-based feature embedding to reconstruct the 4-D LF from 2-D CMs. Specifically, based on the intrinsic linear property of the observation model, we propose a cycle-consistent reconstruction network (CR-Net), which reconstructs the LF in a progressive manner through gradually eliminating the residuals between the back-projected CMs from the reconstructed LF and input CMs.
@@
Table 1: Summary of feature extraction strategies of associated deep learning-based methods, including LF super-resolution, compressive reconstruction, and LF denoising.
@@
LF Super-resolution Yoon et al. (2015) 2-D Conv + SAI stacks No
Wu et al. (2017) 2-D Conv + EPI stacks No
@@
Yeung et al. (2018b) SAS Conv Yes
Jin et al. (2020a) SAS Conv Yes
@@
Wang et al. (2023) 2-D Conv + SAI stacks No
@@ 2.1 Feature Embedding of 4-D LFs
Effective and efficient feature embedding constitutes a fundamental module of diverse learning-based LF tasks and bears a direct relationship to the ultimate performance of the model. Table 1 enumerates the feature extraction techniques utilized in various associated LF tasks, thereby facilitating an intuitive comprehension of the evolution of feature extraction methodologies. Some methods process the 4-D LF from its low-dimensional representation, i.e., 2-D SAI Yoon et al. (2015); Chen et al. (2018a); Wang et al. (2023), 2-D CMs Inagaki et al. (2018); Vadathya et al. (2019), 2-D EPI Wu et al. (2017), and 3-D EPI Wang et al. (2018) volume, which cannot comprehensively explore the distribution of the high-dimensional data. To comprehensively explore the characteristics of high-dimensional LF, Gupta et al. (2017) employed 4-D convolution and MLP to process the 4-D data. Additionally, Yang et al. (2023) devised LF-ResBlock, which is made up of 4-D convolution and deconvolution, to facilitate angular super-resolution. In addition to this simple idea, Yeung et al. (2018b) proposed the SAS convolutional layer that can efficiently process the 4-D LF by sequentially conducting 2-D spatial and angular convolutions. This SAS convolutional layer was employed in several LF tasks, such as angular super-resolution Jin et al. (2020a), compressive reconstruction Guo et al. (2020), and LF denoising Guo et al. (2021). Chen et al. (2022) extended the SAS convolutional to parallel spatial-angular integration blocks (PSAIBs), achieving dense interaction between the spatial and angular domain in a two-stream fashion. Van Duong et al. (2023) employed a hybrid feature extraction module (HFEM) that operates on all three 2-D subspaces (i.e., spatial, angular, and EPI) of the LF image. However, the empirical combination of spatial, angular, and EPI features limits the quality of LF reconstruction.
@@ 2.2 Coded Aperture-based 4-D LF Reconstruction
Recently, some deep learning-based LF denoising methods were proposed, which denoise the LF by training with a large amount of LF data. Chen et al. (2018a) proposed a deep learning-based framework for LF denoising with two sequential CNNs. The first network creates the structural parallax details, and the second restores the view-dependent local energies. In Guo et al. (2021), their deep spatial-angular regularization framework was extended to LF denoising, which implicitly and comprehensively explores the signal distribution based on the observation model of noisy LF images. These learning-based methods have demonstrated excellent performance for light field denoising.
@@ 3 Proposed Probabilistic-based Feature Embedding
To comprehensively explore the spatial-angular information of 4-D LFs, high-dimensional convolution is an intuitive choice that has demonstrated its effectiveness Wang et al. (2018). However, compared to 2-D convolution, it significantly increases the number of parameters, which may cause overfitting and consume significant computational resources. By analogy with the approximation of a high-dimensional filter with multiple low-dimensional filters in the field of signal processing, some researchers have proposed to apply convolutions separately on the spatial Yoon et al. (2015), angular Yeung et al. (2018a), or EPI Wu et al. (2017) domains. However, it is laborious and inconvenient to design an optimal LF feature embedding module manually. Furthermore, the aggregations of different layers also play a crucial role during LF reconstruction. Inspired by the current success of neural architectural search Liu et al. (2018); Wang et al. (2020b), we introduce an adaptive probabilistic-based feature embedding module that builds a probabilistic model to locally choose the LF feature embedding patterns and globally optimize the aggregation patterns of different layers. Thus, both the network architecture and corresponding weights are optimized to achieve efficient and effective LF reconstruction.
@@
To solve Eq. (1), we first approximate the posterior distribution $\mathcal{Q}\left(\mathcal{W}|\mathbb{D}\right)$. According to Gal and Ghahramani (2016), embedding dropout into CNNs could objectively minimize the Kullback–Leibler divergence between an approximation distribution and posterior deep Gaussian process Damianou and Lawrence (2013). Thus, we could approximate the posterior distribution via training a template network with binary masks which follow learnable independent Bernoulli distributions (we name the network with all $\mathbf{U}^{(j)}$ and $\mathbf{V}^{(j)}$ set to $\mathbf{1}$ as the ”template network”). As shown in Fig. 2, both the path for feature aggregation ($\mathbf{U}^{(j)}$) and local feature embedding pattern ($\mathbf{V}^{(j)}$) are replaced by masks of logits $\epsilon\sim\mathcal{B}(p)$, where $\mathcal{B}(p)$ denotes the Bernoulli distribution with probability $p$. However, the classical sampling process is hard to manage a differentiable linkage between the sampling results and the probability. Besides, such dense aggregations in CNNs result in a huge number of feature embeddings. Thus, we also need to explore an efficient and effective way of aggregating those features masked by binary logits. In what follows, we detailedly discuss those two aspects.
@@
where $\widetilde{\mathbf{H}}^{(j)}~{}~{}(1\leq j\leq J)$ denotes the aggregated feature which would be fed into $\mathcal{F}^{(j)}$; $\mathbf{H}^{(k)}~{}~{}(1\leq k\leq j-1)$ represents the feature from the ${k}$-th embedding unit $\mathcal{F}^{(k)}$; $\mathbf{H}^{(0)}$ denotes an LF embedding extracted from the input of $\mathcal{D}^{(t)}(\cdot)$ by a single linear convolutional layer; $\widetilde{\mathbf{U}}^{(j)}(k)$ indicates the ${k}$-th element of the vector $\widetilde{\mathbf{U}}^{(j)}$, which is kept in a range of $[0,1]$ according to its meaning of the sampling probability of $\mathbf{U}^{(j)}$; and $\mathcal{C}_{1\times 1}(\cdot)$ represents a $1\times 1$ kernel to compress the feature embedding and activate them with $ReLU$.
@@
where $\mathcal{C}_{l}^{(j)}(\cdot)~{}~{}(l\in\{1,2,3,4\})$ indicates the previously designed one of the four convolution patterns in $\mathcal{F}^{(j)}$, which are $\mathcal{C}_{spa}^{(j)}(\cdot)$, $\mathcal{C}_{ang}^{(j)}(\cdot)$, $\mathcal{C}_{epih}^{(j)}(\cdot)$, and $\mathcal{C}_{epiv}^{(j)}(\cdot)$, respectively. After applying a $1\times 1$ convolutional layer to compress the embedding, we adopt a further spatial convolution after fusing the feature from different patterns, due to that spatial convolution plays an essential role during feature embedding process.
@@ 4 Proposed Coded Aperture-based 4-D LF Reconstruction
Figure 4: Flowchart of the proposed 4-D LF denoising method.
@@ 6.1 Experiment Settings
Network Settings. We set $T~{}~{}=6$ and $J~{}~{}=8$ for CR-Net and $T~{}~{}=2$ and $J~{}~{}=6$ for our denoising network (see Sections 6.2.3 and 6.3.4 for the ablation study on the effect of $T$ and $J$). In each $\mathcal{D}^{(t)}(\cdot)~{}~{}(0\leq t\leq T-1)$, the kernel size of four types of convolution methods was set to $3\times 3$, and the output feature channel was set to $32$ and $64$ for CR-Net and the denoising network, respectively. Since $\mathcal{D}^{(t)}(\cdot)$ processes residual information, we removed all biases from the convolutional layers. It’s worth noting that all projection layers $\mathcal{P}(\cdot)$ shared the same weights. Additionally, due to their physical interpretation, all parameters in $\mathcal{P}(\cdot)$ were clipped to the range of $[0,1]$ during the training process.
@@
Figure 8: Visual comparisons of the depth maps estimated from the ground-truth LFs and the reconstructed LFs by different methods under task $S=2$ over the HCI+Inria dataset.

### v3

@@ Abstract
The high-dimensional nature of the 4-D light field (LF) poses great challenges in achieving efficient and effective feature embedding, that severely impacts the performance of downstream tasks. To tackle this crucial issue, in contrast to existing methods with empirically-designed architectures, we propose a probabilistic-based feature embedding (PFE), which learns a feature embedding architecture by assembling various low-dimensional convolution patterns in a probability space for fully capturing spatial-angular information. Building upon the proposed PFE, we then leverage the intrinsic linear imaging model of the coded aperture camera to construct a cycle-consistent 4-D LF reconstruction network from coded measurements. Moreover, we incorporate PFE into an iterative optimization framework for 4-D LF denoising. Our extensive experiments demonstrate the significant superiority of our methods on both real-world and synthetic 4-D LF images, both quantitatively and qualitatively, when compared with state-of-the-art methods. The source code will be publicly available at https://github.com/lyuxianqiang/LFCA-CR-NET.
@@ 1 Introduction
To process the 4-D LF, previous learning-based methods simply apply 2-D or 3-D convolutional filters to the stack of sub-aperture images (SAIs) Inagaki et al. (2018); Vadathya et al. (2020); Shin et al. (2018); Zhang et al. (2019), which has a limited ability to explore the angular relations among SAIs. Some advanced network architectures have been designed to model the LF structure better, such as the 4-D convolutional layer Yeung et al. (2019), the epipolar plane images (EPIs) convolutional layer Wu et al. (2017, 2019); Heber et al. (2017), the spatial-angular separable (SAS) convolutional layer Yeung et al. (2018); Jin et al. (2020a); Guo et al. (2020); Jin et al. (2022), the spatial-angular interactive network Wang et al. (2020), the angular deformable convolution Wang et al. (2021a), and the disentangle convolution on macro-pixel image Wang et al. (2023a). However, as these methods combine the spatial and angular features of the LF data in an intuitive or empirical manner, the optimal strategy to model the LF data still needs to be studied. Inspired by the recent advances of Transformer, several Transformer-based methods Wang et al. (2022); Liang et al. (2022, 2023) have been proposed for many LF processing tasks, which can incorporate the information from all angular views and capture long-range spatial dependencies in each SAI.
@@
In this paper, to embed 4-D LF data efficiently and effectively, we propose a probabilistic-based feature embedding module. As shown in Fig. 1, we formulate the problem in a probability space and propose to approximate a maximum posterior distribution (MAP) of a set of carefully defined LF processing events, including both layer-wise spatial-angular feature extraction and network-level feature aggregation. Through droppath from a densely-connected template network, we derive an adaptively probabilistic-based feature embedding, which is sharply contrasted with existing manners that combine spatial and angular features empirically. Building upon the proposed probabilistic-based feature embedding method, we propose two distinct frameworks for the essential tasks of compressive LF imaging and LF image denoising. These tasks are critical in the context of LF data transmission and preprocessing for subsequent applications.
@@
Earlier 4-D LF capturing devices, such as camera array Wilburn et al. (2005) and camera gantry Vaibhav and Andrew , are bulky and costly to capture dense LFs. Although more recent commercial LF cameras, such as Lytro (2016) and RayTrix , are more convenient and efficient, the limited sensor resolution results in a trade-off between spatial and angular resolution. Various manners have been proposed to obtain high-quality 4-D LFs (i.e., 4-D LFs with both high spatial and angular resolution), such as spatial super-resolution Wang et al. (2021a); Jin et al. (2020a); Wang et al. (2023a); Van Duong et al. (2023); Wang et al. (2023b), angular super-resolution (or view synthesis) Jin et al. (2022); Chen et al. (2022); Yang et al. (2023), hybrid lenses Wang et al. (2017); Jin et al. (2023), and coded aperture imaging Inagaki et al. (2018); Vadathya et al. (2020); Guo et al. (2022). Particularly, coded aperture imaging, which encodes the 4-D LF to 2-D coded measurements (CMs) without losing the spatial resolution, demonstrates to be a promising way for high-quality 4-D LF acquisition Wilburn et al. (2005). However, the bottleneck of this manner lies in the ability of the subsequent reconstruction algorithm, which reconstructs the 4-D LF image from the 2-D CMs. Although existing deep learning-based 4-D LF reconstruction methods from CMs Inagaki et al. (2018); Vadathya et al. (2020) have presented significantly better reconstruction quality than conventional ones Babacan et al. (2012); Marwah et al. (2013), they generally employ plain convolutional networks that are purely data-driven without taking the observation model into account. More recently, Guo et al. (2020) proposed an unrolling-based method to link the observation model of coded apertures and deep learning elegantly and improve the LF reconstruction quality significantly. In this paper, we propose a physically interpretable framework incorporated with the probabilistic-based feature embedding to reconstruct the 4-D LF from 2-D CMs. Specifically, based on the intrinsic linear property of the observation model, we propose a cycle-consistent reconstruction network (CR-Net), which reconstructs the LF in a progressive manner by gradually eliminating the residuals between the back-projected CMs from the reconstructed LF and input CMs.
@@
Table 1: Summary of feature extraction strategies used for learning-based LF spatial super-resolution, LF angular super-resolution (or view synthesis), coded aperture-based compressive imaging for LF reconstruction, and LF denoising.
@@
Spatial Super-resolution Yoon et al. (2015) 2-D Conv + SAI stacks No
Chen et al. (2022) 4-D EPI + PSAIB Yes
Wang et al. (2020) SAS Interaction Yes
Wang et al. (2021a) Angular Deformable Conv Yes
Wang et al. (2023a) 2-D Disentangling Conv Yes
Van Duong et al. (2023) 2-D Conv + HFEM Yes
Wang et al. (2022) SAS + Transformer Yes
Liang et al. (2022) SAS + Transformer Yes
Liang et al. (2023) SA Correlation + Transformer Yes
Angular Super-resolution Wu et al. (2017) 2-D Conv + EPI stacks No
@@
Yeung et al. (2018) SAS Conv Yes
Jin et al. (2020b) SAS Conv Yes
Jin et al. (2022) SAS Conv Yes
@@
Wang et al. (2023a) 2-D Disentangling Conv Yes
Van Duong et al. (2023) 2-D Conv + HFEM Yes
@@
Wang et al. (2023c) 2-D Conv + SAI stacks No
@@ 2.1 Feature Embedding of 4-D LFs
Effective and efficient feature embedding constitutes a fundamental module of diverse learning-based LF tasks and bears a direct relationship to the ultimate performance of the model. Table 1 enumerates the feature extraction techniques utilized in various associated LF tasks, thereby facilitating an intuitive comprehension of the evolution of feature extraction methodologies. Some methods process the 4-D LF from its low-dimensional representation, i.e., 2-D SAI Yoon et al. (2015); Chen et al. (2018a); Wang et al. (2023c), 2-D CMs Inagaki et al. (2018); Vadathya et al. (2020), 2-D EPI Wu et al. (2017), and 3-D EPI Wang et al. (2018) volume, which cannot comprehensively explore the distribution of the high-dimensional data. To comprehensively explore the characteristics of high-dimensional LF, Gupta et al. (2017) employed 4-D convolution and MLP to process the 4-D data. Additionally, Yang et al. (2023) devised LF-ResBlock, which is made up of 4-D convolution and deconvolution, to facilitate angular super-resolution. In addition to this simple idea, Yeung et al. (2018) proposed the SAS convolutional layer that can efficiently process the 4-D LF by sequentially conducting 2-D spatial and angular convolutions. This SAS convolutional layer was employed in several LF tasks, such as angular super-resolution Jin et al. (2022), compressive reconstruction Guo et al. (2020), and LF denoising Guo et al. (2022). Chen et al. (2022) extended the SAS convolutional to parallel spatial-angular integration blocks (PSAIBs), achieving dense interaction between the spatial and angular domain in a two-stream fashion. Wang et al. (2023a) proposed a disentangling mechanism in the 2-D macro-pixel image to extract spatial, angular, and EPI features. However, the empirical combination of spatial, angular, and EPI features may limit the quality of LF reconstruction. Thus, several methods are proposed to combine the high-dimensional feature in some smarter ways Lyu et al. (2021); Van Duong et al. (2023). For example, Van Duong et al. (2023) employed a hybrid feature extraction module (HFEM) that operates on all three 2-D subspaces (i.e., spatial, angular, and EPI) of the LF image. Besides these CNN-based methods, many Transformer-based methods Wang et al. (2022); Liang et al. (2022, 2023) have been proposed for different LF processing tasks.
@@ 2.2 Coded Aperture-based 4-D LF Reconstruction
Recently, some deep learning-based LF denoising methods were proposed, which denoise the LF by training with a large amount of LF data. Chen et al. (2018a) proposed a deep learning-based framework for LF denoising with two sequential CNNs. The first network creates the structural parallax details, and the second restores the view-dependent local energies. In Guo et al. (2022), their deep spatial-angular regularization framework was extended to LF denoising, which implicitly and comprehensively explores the signal distribution based on the observation model of noisy LF images. Wang et al. (2023c) utilized the multiple stream progressive restoration network framework and averaged view stacks to LF denoising. These learning-based methods have demonstrated excellent performance for light field denoising.
@@ 3 Proposed Probabilistic-based Feature Embedding
To comprehensively explore the spatial-angular information of 4-D LFs, high-dimensional convolution is an intuitive choice that has demonstrated its effectiveness Wang et al. (2018). However, compared to 2-D convolution, it significantly increases the number of parameters, which may cause overfitting and consume significant computational resources. By analogy with the approximation of a high-dimensional filter with multiple low-dimensional filters in the field of signal processing, some researchers have proposed to apply convolutions separately on the spatial Yoon et al. (2015), angular Yeung et al. (2019), or EPI Wu et al. (2017) domains. However, it is laborious and inconvenient to design an optimal LF feature embedding module manually. Furthermore, the aggregations of different layers also play a crucial role during LF reconstruction. Inspired by the current success of neural architectural search Liu et al. (2019); Wang et al. (2021b), we introduce an adaptive probabilistic-based feature embedding module that builds a probabilistic model to locally choose the LF feature embedding patterns and globally optimize the aggregation patterns of different layers. Thus, both the network architecture and corresponding weights are optimized to achieve efficient and effective LF reconstruction.
@@
To solve Eq. (1), we first approximate the posterior distribution $\mathcal{Q}\left(\mathcal{W}|\mathbb{D}\right)$. According to Gal and Ghahramani (2016), embedding dropout into CNNs could objectively minimize the Kullback–Leibler divergence between an approximation distribution and posterior deep Gaussian process Damianou and Lawrence (2013). Thus, we could approximate the posterior distribution via training a template network with binary masks that follow learnable independent Bernoulli distributions (we name the network with all $\mathbf{U}^{(j)}$ and $\mathbf{V}^{(j)}$ set to $\mathbf{1}$ as the template network). As shown in Fig. 2, both the path for feature aggregation ($\mathbf{U}^{(j)}$) and local feature embedding pattern ($\mathbf{V}^{(j)}$) are replaced by masks of logits $\epsilon\sim\mathcal{B}(p)$, where $\mathcal{B}(p)$ denotes the Bernoulli distribution with probability $p$. However, the classical sampling process makes it hard to manage a differentiable linkage between the sampling results and the probability. Besides, such dense aggregations in CNNs result in a huge number of feature embeddings. Thus, we also need to explore an efficient and effective way of aggregating those features masked by binary logits. In what follows, we detailedly discuss those two aspects.
@@
where $\widetilde{\mathbf{H}}^{(j)}~{}~{}(1\leq j\leq J)$ denotes the aggregated feature which would be fed into $\mathcal{F}^{(j)}$; $\mathbf{H}^{(k)}~{}~{}(1\leq k\leq j-1)$ represents the feature from the ${k}$-th embedding unit $\mathcal{F}^{(k)}$; $\mathbf{H}^{(0)}$ denotes an LF embedding extracted from the input of $\mathcal{D}^{(t)}(\cdot)$ by a single linear convolutional layer; $\widetilde{\mathbf{U}}^{(j)}(k)$ indicates the ${k}$-th element of the vector $\widetilde{\mathbf{U}}^{(j)}$, which is random initialized and kept in a range of $[0,1]$ according to its meaning of the sampling probability of $\mathbf{U}^{(j)}$; and $\mathcal{C}_{1\times 1}(\cdot)$ represents a $1\times 1$ kernel to compress the feature embedding and activate them with $ReLU$.
@@
where $\mathcal{C}_{l}^{(j)}(\cdot)~{}~{}(l\in\{1,2,3,4\})$ indicates the previously designed one of the four convolution patterns in $\mathcal{F}^{(j)}$, which are $\mathcal{C}_{spa}^{(j)}(\cdot)$, $\mathcal{C}_{ang}^{(j)}(\cdot)$, $\mathcal{C}_{epih}^{(j)}(\cdot)$, and $\mathcal{C}_{epiv}^{(j)}(\cdot)$, respectively. After applying a $1\times 1$ convolutional layer to compress the embedding, we adopt a further spatial convolution after fusing the feature from different patterns, due to that spatial convolution plays an essential role during the feature embedding process.
@@ 4 Proposed Coded Aperture-based 4-D LF Reconstruction
Figure 4: Illustration of the process of the proposed 4-D LF denoising method, namely DN-Net.
@@ 6.1 Experiment Settings
Network Settings. We set $T~{}~{}=6$ and $J~{}~{}=8$ for CR-Net and $T~{}~{}=2$ and $J~{}~{}=6$ for our denoising network (see Sections 6.2.3 and 6.3.5 for the ablation study on the effect of $T$ and $J$). In each $\mathcal{D}^{(t)}(\cdot)~{}~{}(0\leq t\leq T-1)$, the kernel size of four types of convolution methods was set to $3\times 3$, and the output feature channel was set to $32$ and $64$ for CR-Net and DN-Net, respectively. Since $\mathcal{D}^{(t)}(\cdot)$ processes residual information, we removed all biases from the convolutional layers. It’s worth noting that all projection layers $\mathcal{P}(\cdot)$ shared the same weights. Additionally, due to their physical interpretation, all parameters in $\mathcal{P}(\cdot)$ were clipped to the range of $[0,1]$ during the training process.
@@
Figure 8: Visual comparisons of ground-truth depth maps and the estimated depth maps from the reconstructed LFs by different methods under task $S=2$ over the HCI+Inria dataset.